OpenStudy (anonymous):

a great circle on a standard bowling ball is about 26 inches. To the nearest unit, how many square inches is the surface area of 100 bowling balls? LOL Idk this

OpenStudy (anonymous):

the surface area of a sphere is \[4\pi r^2\] to solve for the radius, we know that the equation for the circumference of a circle is \[2\pi r\] and the circumference of each bowling ball is 26 inches, which means the radius of each bowling wall is 26/(2*pi) ----> 13/pi = r so now plug the radius into the equation for the surface area of a sphere, except we need to solve for the surface area for 100 bowling balls, so we multiply it by 100 100*4*pi*(13/pi)^2 = surface area of 100 bowling balls